# Diagnosing Quantum Many-body Chaos in Non-Hermitian Quantum Spin Chain via Krylov Complexity

### v1

@@ Abstract
Conclusions.— In conclusion, we have studied non-Hermitian quantum chaos in the disordered XY model, revealing two critical phase transitions including the chaotic-non-chaotic transition and Krylov localization. Through examing the Krylov complexity and Lanczos coefficients characterized by the Krylov variance and reciprocity, we identify the critical points and exponents of these phase transitions. Our results are corroborated by established measures including entanglement entropy and complex level spacings, which showcase the applicability of Krylov complexity and the associated measures in the study of quantum chaos. The sign-flip of Krylov reciprocity provides a quantitatively more precise criterion for chaos suppression than the entanglement entropy. Furthermore, we observe intricate chaotic behavior in non-Hermitian systems, where prethermalization may exist in the Krylov delocalized regime.
The non-Hermitian spin model can be realized in Rydberg atom arrays, where spin states are encoded by two Rydberg states [50]. The XY interactions correspond to the dipole-dipole interactions between the Rydberg states, and the dissipation can be induced by coupling the Rydberg states to low-lying electronic states [51]. Randomness of the local fields can be introduced by AC-Stark shifts with addressing lasers. In generic quantum simulators, non-Hermiticity can be realized through Sz.-Nagy dilation [52, 53] and linear combination of unitaries [54, 55]. Finally, to measure the Krylov complexity, it has been proposed to apply measurable bases for estimation [56], and the measure of Loschmidt echo is potentially related to the ergodicity breaking [Fig. 1(d)].
Acknowledgements.
We are grateful to Dong-Ling Deng, Dong Yuan, Si Jiang, Chang Liu (SQZI), Chang Liu (NUS), Tianyi Yan, Xiaopeng Li, Xingze Qiu and Saud Čindrak for fruitful discussions. Y.Z. acknowledges support from the Shanghai Qi Zhi Institute Innovation Program SQZ202317. W.X. acknowledges support from the National Natural Science Foundation of China/Hong Kong RGC Collaborative Research Scheme (Project CRS CUHK401/22) and New Cornerstone Science Foundation. L.L. acknowledges support from the National Key Research and Development Program of China (Grant No. 2021YFA1402003) and the National Natural Science Foundation of China (Grant Nos. 12374329 and U21A6006). W.L. acknowledges support from the EPSRC through Grant No. EP/W015641/1, the Going Global Partnerships Programme of the British Council (Contract No. IND/CONT/G/22-23/26), and the International Research Collaboration Fund of the University of Nottingham.

### v2

@@ Appendix D Fitting of Lanczos coefficients with Tsallis $q$-log statistics
Thus, the assumption $b_{n}\propto\sqrt{1-n/s^{L}}$ in Ref. [26] can be regarded as a special case with $q=0$. We fit the average Lanczos coefficients $\langle|b_{n}|\rangle$ and $\langle|c_{n}|\rangle$ using the form
$\langle|b_{n}|\rangle,\langle|c_{n}|\rangle\propto\sqrt{1-(n/2^{L})^{1-q}},$ (19)
with $q=0$ and $0.75$, as shown in Fig. S5. We find that $q=0.75$ provides a better approximation for the values of Lanczos coefficients in the regime where $n>L$, as discussed in the main text.
Appendix E Krylov reciprocity
In order to show the $d$-dependence of the Krylov reciprocity, we show the values of $R_{K}^{(d)}$ for $d=3$ to $8$ in Fig. S6(a1-a6). The crossing behavior is universal for all the choices of $d$, while their finite-size scaling has some small differences as shown in Fig. S6(b,c).
Figure S6: (a1-a6) Krylov reciprocity $R_{K}^{(d)}$ for $d=3$ to 88. The insets are scaling collapse results for $x=(W_{\gamma}-W_{C})L^{\alpha}$ and $y=R_{K}^{(d)}$. The critical points $W_{C}$ (b) and exponents $\alpha$ (c) are shown to have a shared plateau for $d=4$ to $6$, which are selected for identification of phase transition.
Here, we note that the finite-size collapse is carried out by minimizing the cosine similarity of the scaled data with the Nelder-Mead algorithm. This method is also applied to other finite-scaling analysis throughout this work.
We find that for $d=3$ to $7$, the collapsed critical point, $W_{C}$, exhibits a plateau, and there is a plateau for $d=4$ to $6$ for the critical exponent, $\alpha$. Taking the average values for $d=4$ to $6$, we find that $\langle W_{C}\rangle=1.647\pm 0.014$, and $\langle\alpha\rangle=0.820\pm 0.003$. The relative mean deviations are $0.85\%$ and $0.36\%$ for $\langle W_{C}\rangle$ and $\langle\alpha\rangle$, respectively.
Appendix F Bipartite entanglement entropy
In this section, we present the bipartite entanglement entropy of each eigenstate, as depicted in Fig. S7. The entanglement is quantified using the von Neumann entropy for half of the chain, $S=-\textrm{Tr}\left(\rho_{L/2}\ln\rho_{L/2}\right)$, where the reduced density matrix $\rho_{L/2}$ is obtained by tracing out the degrees of freedom of the chain’s other half with indices $j>L/2$. The variance of entanglement entropy is taken by deviation of the mean entropy over different random configurations [53].
Figure S7: Bipartite entanglement entropy versus the real part of the eigenvalues of each eigenstate. The orange lines are the Page value. Here we set $L=12$ and $S_{\rm Page}=6\ln 2-1/2\approx 3.659$.
Generally, in the chaotic regime, the bipartite entanglement entropy scales extensively with subsystem volume, which is half of the total system size, $L/2$. The maximum entropy approaches the Page value, $S_{\rm Page}=(L\ln 2-1)/2$ [53]. As disorder strength $W_{\gamma}$ increases and the system transitions to the nonchaotic phase, the entanglement entropy of eigenstates decreases and their distribution narrows significantly. This compression shifts the shape of the distribution from a broad arc to a cigar-like profile. At extreme disorder (e.g., $W_{\gamma}=4.0$), the entropy approaches zero, signaling eigenstates that are nearly product states. This near-factorization of wavefunctions strongly suggests the emergence of integrability in the strongly dissipative limit.
In the chaotic regimes where $W_{\gamma}$ is small, the mean value of the entanglement entropy, $\bar{S}$, increases as shown in Fig. S8(a). At this point, no discernible difference is observed between Hermitian and non-Hermitian chaos, as $S$ approaches constant values when $W_{\gamma}\to 0$. As $W_{\gamma}$ increases, $\bar{S}$ tends to diminish, which is consistent with the hypothesis that the system’s wave function evolves into product states. Additionally, we have conducted a finite-size scaling analysis, as shown in Fig. S8(b). The critical value $W^{*}$ as well as the exponents $\alpha$ and $\beta$ are obtained by finite-size scaling and $W^{*}=3.236$ is in proximity to the transition point where the characteristic complex level spacing $\langle r\rangle$ descends to $2/3$, indicative of a 2D Poisson distribution.
Figure S8: (a) Bipartite entanglement entropy for varying system size $L$ and non-Hermitian disorder $W_{\gamma}$. The red dashed line gives $W^{*}$ obtained from the finite-size scaling. (b) Finite-size scaling of $S$. The Red line is the collapsed universal function by averaging data within the interval for optimization.
Conversely, the examination of the standard deviation of the bipartite entanglement entropy, $\sigma_{S}$, along with Krylov reciprocity, indicates a phase transition at $W_{C}\approx 1.763$ and $1.647$, respectively, as it departs from the chaotic regime of the AI class where $\langle r\rangle_{\textrm{AI}^{\dagger}}=0.7222$.